OpenStudy (anonymous):

How to solve for x: (X+3)^3=2

OpenStudy (johnweldon1993):

First take the cubed root of both sides (this will eliminate the ^3 on the left side so now you have x + 3 = cubed root 2 to solve for x...just subtract 3 from both sides so \[x = \sqrt[3]{2} - 3\]

OpenStudy (tkhunny):

That's a great was to proceed it you want only one of the three solutions. Use this \((a^{3} - b^{3}) = (a-b)(a^{2}+ab+b^{2})\) Consider that \(b = \sqrt[3]2\)